OpenStudy (anonymous):

1. Write using exponents: (x)(x)(x)(x)(y)(y) 2. Simplify: 25 + 3² • 5 + 4² 3. Simplify: 34 + 2 • 34 4. Simplify: Evaluate: 3x² – 4 for x = 3 5. Find the GCF: 18x³ and 30x5 1. Write using exponents: (x)(x)(x)(x)(y)(y) 2. Simplify: 25 + 3² • 5 + 4² 3. Simplify: 34 + 2 • 34 4. Simplify: Evaluate: 3x² – 4 for x = 3 5. Find the GCF: 18x³ and 30x5 @Writing

OpenStudy (anonymous):

x^4y^2 25 + 9*5 + 16 = 86 34 + 68 = 102 3*9 - 4 = 23 6x^3

OpenStudy (anonymous):

3=102